### v1

# Variations on a theme of empty polytopes

Abstract
Given a set $S\subseteq\mathbb{R}^{d}$, an empty polytope has vertices in $S$ but contains no other point of $S$. Empty polytopes are closely related to so-called Helly numbers, which extend Helly’s theorem to more general point sets in $\mathbb{R}^{d}$. We improve bounds on the number of vertices in empty polytopes in exponential lattices, arithmetic congruence sets, and 2-syndetic sets.
We also study hollow polytopes, which have vertices in $S$ and no points of $S$ in their interior. We obtain bounds on the number of vertices in hollow polytopes under certain conditions, such as the vertices being in general position.
Finally, we obtain relatively tight asymptotic bounds for polytopes which do not contain lattice segments of large length.
1 Introduction
Helly’s theorem is a fundamental result in discrete and convex geometry. It says that in a finite collection $\mathcal{C}$ of convex sets in $\mathbb{R}^{d}$, if the intersection of any $d+1$ or fewer sets in $\mathcal{C}$ is nonempty, then $\bigcap\mathcal{C}$ is nonempty, as well. Helly’s theorem was first proven by Eduard Helly in 1913 (though not published until 1923 [10]) and independently by Radon [15] in 1921. Helly’s theorem was the invitation to a veritable cornucopia of results in discrete geometry, and the last century has led to dozens of variations on Helly’s theorem. Bárány and Kalai’s survey [4] provides a recent and extensive look at results in the broad area of related results, while the survey by Amenta, de Loera, and Soberón [2] focuses more specifically on Helly’s theorem.
In 1973, Doignon proved a Helly-type theorem for the integer lattice [8].
Theorem (Doignon’s theorem).
Let $\mathcal{C}$ be a finite collection of convex sets in $\mathbb{R}^{d}$. If the intersection of any $2^{d}$ or fewer sets in $\mathcal{C}$ contains an integer point, then $\bigcap\mathcal{C}$ does, too.
This theorem was also independently discovered by Bell [6] and Scarf [17]. A few years later, Hoffman extended Bell’s result by connecting Helly-type theorems with polytopes.
Given a set $S\subseteq\mathbb{R}^{d},$ the Helly number of $S,$ denoted $h(S),$ is the smallest $h$ such that the following Helly-type theorem holds:
Let $\mathcal{F}$ be a finite family of convex sets in $\mathbb{R}^{d}.$ If every $h$ or fewer sets in $\mathcal{F}$ contains a point of $S$ in their intersection, then the intersection of all sets in $\mathcal{F}$ contains a point in $S.$
If no such $h$ exists, we say $h(S)=\infty.$ Helly’s theorem says that $h(\mathbb{R}^{d})\leq d+1$ (while the collection of facets of a simplex shows that $h(\mathbb{R}^{d})\geq d+1)$, and Doignon’s theorem says that $h(\mathbb{Z}^{d})\leq 2^{d}$.
Moreover, a subset $T\subseteq S$ is called empty in $S$ if $\operatorname{conv}(T)\cap S$ only contains vertices of $\operatorname{conv}(T)$. We call a set $S$ discrete if it has no accumulation points.
Theorem (Hoffman [11]).
If $S\in\mathbb{R}^{d}$ and $S$ is discrete, then $h(S)$ is equal to the maximum size of an empty subset of $S$.
Hoffman actually proved a version of this result that holds for any subset $S\subseteq\mathbb{R}^{d}$, but we won’t need that more general result. Since the unit hypercube $\{0,1\}^{d}$ is empty in $\mathbb{Z}^{d}$, we have $h(\mathbb{Z}^{d})\geq 2^{d}$; this shows that Doignon’s theorem is tight. Further works on Helly numbers have considered mixed integer subsets $\mathbb{Z}^{r}\times\mathbb{R}^{d-r}$ [3], algebraic sets [2], or unions of lattices [9].
In the first part of this paper, we study Helly numbers of product sets, or sets of the form $S=A^{d}$, for some $A\subseteq\mathbb{R}$. Dillon [7] proved a general criterion on product sets which implies in particular that if $p$ is a polynomial with degree at least $2$ and $A=\{p(n):n\in\mathbb{Z}\},$ then $h(A^{2})=\infty$ for all $d\geq 2.$ Based on this result, one might conjecture that the finiteness of $h(A^{2})$ is related to the sparseness of the set $A$. However, this is ruled out in [7] by the construction of a set $A\subseteq\mathbb{Z}$ whose consecutive elements differ by at most $2$ (a “$2$-syndetic” set) such that the $A^{2}$ has arbitrarily large empty polygons, and thus arbitrarily large Helly number. In Section 2.2, we strengthen this result by constructing a $2$-syndetic set with a single polygon with infinitely many vertices.
Dillon’s method, however, gives no information for exponential lattices: sets of the form $L_{d}(\alpha)=\{\alpha^{n}:n\in\mathbb{N}_{0}\}^{d}$. Ambrus, Balko, Frankl, Jung, and Naszódi [1] studied these sets in two dimensions, and they obtained upper bounds for all $\alpha>1$ and exact values for $\alpha\in[\tfrac{1+\sqrt{5}}{2},\infty).$
Theorem (Ambrus, Balko, Frankl, Jung, Naszódi [1]).
Let $\alpha>1.$
  • $\circ$
    If $\alpha\geq 2,$ then $h(L_{2}(\alpha))=5.$
  • $\circ$
    If $\alpha\in[\tfrac{1+\sqrt{5}}{2},2),$ then $h(L_{2}(\alpha))=7.$
  • $\circ$
    If $\alpha\in(1,\tfrac{1+\sqrt{5}}{2}),$ then $h(L_{2}(\alpha))\leq 3\lceil\log_{\alpha}(\tfrac{\alpha}{\alpha-1})\rceil+3.$
In Section 2.1.1, we obtain a stronger bound, with an approach that is shorter and more geometric:
Theorem 1.1.
For $\alpha>1,$
$h(L_{2}(\alpha))\leq 2\left\lceil\log_{\alpha}\left(\frac{\alpha}{\alpha-1}\right)\right\rceil+3.$
This method recovers the exact values of $h(L_{2}(\alpha))$ for $\alpha>\tfrac{1+\sqrt{5}}{2}.$ It also allows us to fully characterize all maximal empty polygons when $\alpha\geq 2.$
Ambrus et al. [1] also proved a lower bound for $h(L_{2}(\alpha))$ by constructing an empty polygon with vertices on a hyperbola. They showed that for $\alpha>1,$ we have $h(L_{2}(\alpha))\geq\sqrt{\frac{1}{\alpha-1}}$. In Section 2.1.2, we extend this bound to all dimensions:
Theorem 1.2.
For $\alpha>1$ and $d>1,$
$h(L_{d}(\alpha))\geq\binom{k+d-1}{d-1},\vspace{-0.25\baselineskip}$
where $k=\left\lfloor\sqrt{\tfrac{1}{\alpha-1}}\,\right\rfloor.$
In Section 2.3, we investigate Helly numbers of powers of arithmetic congruence sets, which are sets of the form $A=S+m\mathbb{Z}$ for some subset $S\subseteq\{0,1,2\dots,m\}$. Garber [9] showed that $h(S)\leq k2^{d}$ if $S$ is the union of $k$ translates of $\mathbb{Z}^{d}$; if $d=2$, he proved that $h(S)\leq k+6$. If $A=S+m\mathbb{Z}$, then $A^{d}$ is the union of $|S|^{d}$ translates of $\mathbb{Z}^{d}$, so $h(A^{2})\leq|S|^{2}+6$ and $h(A^{d})\leq\big{(}2|S|\big{)}^{d}$. We improve these bounds for certain arithmetic congruence sets.
Theorem 1.3.
Let $m$ be a positive integer with smallest prime factor $p$ and let $k$ and $d$ be positive integers such that $d<\tfrac{p(m-1)}{m(m-k)}.$ If $A=S+m\mathbb{Z}$ where $|S|=k,$ then
$h(A^{d})\leq k^{d}.$
This theorem is strongest when $m$ is itself prime, in which case the inequality reduces to $d<(m-1)/(m-k)$. The corollary gives a significantly improved bound in this case:
Corollary 1.4.
For any prime $p$ and any $\alpha>1-\frac{1}{d}\big{(}1-\frac{1}{p}\big{)}$, and for any set $S\subseteq\{1,2,\dots,p\}$ with $|S|\geq\alpha p$, define $A=S+p\mathbb{Z}$. In this situation,
$h(A^{d})\leq|S|^{d}.$
In the second part of this paper, we investigate related questions for variations of the “empty polytope” property. There is a large literature on so-called hollow polytopes, which are integer polytopes that contain no lattice point in their interior. (See [12] for a recent survey.) Much of the focus is on classification of hollow polytopes up to unimodular equivalence, and there is a strong connection with the celebrated Flatness Theorem (see Theorem 3.4 for a statement of the theorem). Our focus instead is on bounding the number of vertices of a hollow polytope, and we prove the following results:
Theorem 1.5.
If $P$ is a hollow polytope in $\mathbb{Z}^{d}$ in which every facet is a simplex, then $P$ has at most $2^{d}$ vertices. Furthermore, if the vertices of $P$ are in general position, then $P$ has $O\big{(}d^{2}(\log d)^{3}\big{)}$ vertices.
As we describe in Section 3.1, to obtain any bound on the number of vertices, some restriction (such as simplicial facets or vertices general position) must be imposed; these two conditions seem fairly reasonable.
Finally, we prove bounds on the number of vertices of polytopes that contain no “lattice segments” of large length.
Theorem 1.6.
If $P$ is a convex lattice polytope in $\mathbb{R}^{d}$ which does not contain a set of the form $\{x,x+y,x+2y,\dots,x+ky\}$ with $x,y\in\mathbb{Z}^{d}$ and $y\neq 0$, then $P$ has $O_{d}(k^{d-1})$ vertices.
We also construct a polytope with $\Theta_{d}(k^{(d-1)\frac{d}{d+1}})$ vertices containing no such set.
We conclude in Section 4 with several open problems.
2 Helly-type theorems
2.1 Exponential lattices
2.1.1 Upper Bound
Ambrus et al. [1] prove establish an upper bound on $h\big{(}L_{2}(\alpha)\big{)}$ by considering an empty polygon and dividing its edges into four quadrants. Then they bound the number of edges in each quadrant by making use of the emptiness property.
We also consider an empty polygon in an exponential lattice, and in one case (where all the edge slopes are nonnegative), we repeat the analysis in [1]. Our remaining analysis, however, considers a different set of points, which is how we achieve an improved bound with reduced casework.
Lemma 2.1 (Lemma 10 + Corollary 11 in [1]).
An empty polygon in $L_{2}(\alpha)$ that contains only edges with nonnegative slope has at most $22\lceil\log_{\alpha}(\frac{\alpha}{\alpha-1})\rceil+$ 2 edges.
Proof of Theorem 1.1.
Consider an empty polygon $\mathcal{P}$ in $L_{2}(\alpha).$ If all edges of $\mathcal{P}$ have nonnegative slope, then Lemma 2.1 is sufficient.
Otherwise, $\mathcal{P}$ contains vertices $a=(\alpha^{p},\alpha^{q})$ and $b=(\alpha^{r},\alpha^{s})$ such that $p<r$ and $q>s.$ We may choose $a$ and $b$ so that no vertices are above and to the left of $a$ or below and to the right of $b$. A diagram is shown in Figure 1.
Let $c=(\alpha^{r-1},\alpha^{q-1})$ and let $\ell_{1},\ell_{2}$ be the horizontal and vertical lines passing through $c$ respectively. Let $x$ and $y$ be the intersections of $\overline{ab}$ with $\ell_{1}$ and $\ell_{2}$ respectively. We define $i,j\in\mathbb{N}$ such that $d:=(\alpha^{r-1-i},\alpha^{q-1})$ is the rightmost point of $\ell_{1}\cap L_{2}(\alpha)$ strictly to the left of $\overline{ab},$ and $e:=(\alpha^{r-1},\alpha^{q-1-j})$ is the uppermost point of $\ell_{2}\cap L_{2}(\alpha)$ strictly below $\overline{ab}$. It is possible for $c$ to lie below $\overline{ab},$ in which case the points $c,d,e$ coincide; this will not affect our proof.
[…]
Figure 1: The named points in the proof of Theorem 1.1.
To bound $i$ and $j,$ let $a^{\prime}=(0,\alpha^{q})$ and $b^{\prime}=(\alpha^{r},0).$ The lines $\overline{a^{\prime}b^{\prime}}$ and $\ell_{1}$ intersect at $x^{\prime}=(\alpha^{r}-\alpha^{r-1},\alpha^{q-1}).$ Since $x^{\prime}$ is strictly to the left of $x$ and the first coordinate of $x$ is less than $\alpha^{r-i}$, we have $\alpha^{r}-\alpha^{r-1}<\alpha^{r-i}$; in other words,
$i<\log_{\alpha}\left(\frac{\alpha}{\alpha-1}\right).$
The same bound holds for $j$ by considering the intersection of $\overline{a^{\prime}b^{\prime}}$ and $\ell_{2}.$
Our next goal is to deduce restrictions on the possible locations of the vertices of $\mathcal{P}$. Let $f$ be the point such that $cde\!f$ is a rectangle. Let $g$ be the lattice point immediately to the right of $a,$ and let $h$ be the lattice point immediately above $b.$ Consider a point $p$ below $\overline{ab}$ lying outside pentagon $xye\!f\!d.$
  • $\circ$
    If $p$ lies below and to the left of $d,$ then triangle $pab$ contains $d.$
  • $\circ$
    If $p$ lies below and to the left of $e,$ then triangle $pab$ contains $e.$
  • $\circ$
    Otherwise, $p$ lies below and to the right of $b$ or above and to the left of $a,$ violating our assumption.
Therefore, all vertices of $\mathcal{P}$ below $\overline{ab}$ must lie within $xye\!f\!d.$
Similarly, consider a point $p$ above $\overline{ab}$ lying outside triangle $cxy.$
  • $\circ$
    If $p$ lies on or above $\ell_{1},$ then triangle $pab$ contains $g.$
  • $\circ$
    If $p$ lies on or to the the right of $\ell_{2},$ then triangle $pab$ contains $h.$
Thus, all vertices of $\mathcal{P}$ above $\overline{ab}$ must either be $g$ or $h,$ or lie within $cxy.$
To finish, we bound the number of vertices in rectangle $cdf\!e.$ For any given $x$-coordinate, no two vertices with that $x$-coordinate can lie on the same side of $\overline{ab}$ (and similarly for $y$-coordinates). This implies that the number of vertices in rectangle $cd\!f\!e$ is at most
$\min(i,j)+\min(i+1,j+1)\leq 2\left\lceil\log_{\alpha}\left(\frac{\alpha}{\alpha-1}\right)\right\rceil-1.$
Including the vertices $a,b,g,h$ yields the desired bound. ∎
Based on this proof, we can easily characterize all maximal empty polygons for $\alpha\geq 2.$
Corollary 2.2.
If $\alpha\geq 2,$ then all empty pentagons in $L_{2}(\alpha)$ have vertices of the form $(\alpha^{p},\alpha^{q}),(\alpha^{r},\alpha^{s}),(\alpha^{r-1},\alpha^{q-1}),(\alpha^{p+1},\alpha^{q}),(\alpha^{r},\alpha^{s+1}),$ where $p<r$ and $q>s.$
Proof.
Lemma 2.1 implies that any empty polygon in which every edge has nonnegative slope has at most 4 vertices. So the pentagon must contain at least one edge with negative slope; we therefore adopt the terminology of our proof of Theorem 1.1.
If $\alpha\geq 2,$ then $\log_{\alpha}(\tfrac{\alpha}{\alpha-1})\leq 1.$ This means both $i$ and $j$ in the above proof are $0,$ which means $C=D=E.$ So the only possible vertices in an empty polygon are $A,B,C,G,H,$ which are exactly the points stated in the corollary. ∎
2.1.2 Lower Bounds
We now prove a lower bound for $h\big{(}L_{d}(\alpha)\big{)}$.
Proof of Theorem 1.2.
Consider the set of points $X$ in $L_{d}(\alpha)$ which lie on the convex curve $\prod_{i=1}^{d}x_{i}=\alpha^{k}$ We claim that $\operatorname{conv}(X)$ is empty in $L_{d}(\alpha)$. Since $\prod_{i=1}^{d}x_{i}=\alpha^{k}$ is a convex function, the only possible points of $L_{d}(\alpha)$ that lie in $\operatorname{conv}(X)$ satisfy $\prod_{i=1}^{d}x_{i}\geq\alpha^{k}$. All the points satisfying $\prod_{i=1}^{d}x_{i}=\alpha^{k}$ are vertices of $\operatorname{conv}(X)$, so any other point in $\operatorname{conv}(X)\cap L_{d}(\alpha)$ satisfies $\prod_{i=1}^{d}x_{i}\geq\alpha^{k+1}$. On the other hand, the polytope $\operatorname{conv}(X)$ lies completely in the half-space $\sum_{i=1}^{d}x_{i}\leq\alpha^{k}+d-1$. (The bounding hyperplane of this half-space coincides with the facet determined by the vertices $(1,1,\ldots,\alpha^{k}),\ldots,(\alpha^{k},1,\ldots,1)$.) Therefore, to show that $\operatorname{conv}(X)$ is empty in $L_{d}(\alpha)$, it suffices to show that the curve $\prod_{i=1}^{d}x_{i}\geq\alpha^{k+1}$ lies strictly above the hyperplane $\sum_{i=1}^{d}x_{i}=\alpha^{k}+d-1$.
The point on $\prod_{i=1}^{d}x_{i}\geq\alpha^{k+1}$ lying closest to the hyperplane is $(\alpha^{(k+1)/d},\ldots,\alpha^{(k+1)/d})$, so it suffices to show that
$d\alpha^{(k+1)/d}\geq\alpha^{k}+d-1.$
We will apply the substitution $\alpha=1+s^{2}$. Using the inequalities $(1+s^{2})^{(k+1)/d}\geq 1+\tfrac{k+1}{d}s^{2}$ (by the binomial theorem) and $(1+s^{2})^{k}<e^{s^{2}k}$ (by Taylor approximation), we can see that it is enough to prove that
$d\left(1+\frac{k+1}{d}s^{2}\right)\geq e^{s^{2}k}+n-1.$
Our particular choice of $k$ is $k=1/s$, so the previous inequality simplifies to
$s^{2}+s+1\geq e^{s},$
which follows from the Taylor expansion of $e^{x}.$ Thus we have found an empty polytope.
The number of vertices in this polytope is the number of ordered $d$-tuples of nonnegative integers $(n_{1},n_{2},\dots,n_{d})$ with $\prod_{i=1}^{k}\alpha^{n_{i}}=\alpha^{k}$; in other words, the number of $d$-tuples $(n_{1},\dots,n_{d})$ with $\sum_{i=1}^{d}n_{i}=k$. This is well known to be $\binom{k+d-1}{d-1}$ (for example, by the “stars and bars” method). ∎
2.2 Infinite Empty Polygons in Integer Sets
A set $A\subseteq\mathbb{Z}$ is called $2$-sydentic if for every $n\in\mathbb{Z}$, either $n\in A$ or $n+1\in A$ (or both). To construct a $2$-syndetic set $A$ for which $A^{2}$ has infinite Helly number, Dillon [7] constructed a sequence of successively larger empty polygons using a sequence of rational approximation to a fixed irrational number. In [1], Ambrus et al. prove that the “Fibonacci lattice” contains an infinite empty polygon. We combine these approaches to construct a 2-syndetic set containing an empty polygon with infinitely many vertices.
Proposition 2.3.
There is a set $A\subseteq\mathbb{Z}$ such that
  • $\circ$
    for every $n\in\mathbb{Z}$, either $n\in A$ or $n+1\in A$ (or both), and
  • $\circ$
    $A\times A$ contains an infinite empty polygon.
Proof.
Define the Fibonacci numbers by $F_{0}=0,F_{1}=1,$ and $F_{n}=F_{n-1}+F_{n-2}.$ Let $\phi=\tfrac{1+\sqrt{5}}{2}$ and $\psi=\tfrac{1-\sqrt{5}}{2}.$ Let $\mathcal{P}$ be the convex hull of the following points:
$\{p_{i}=(-F_{2i},2F_{2i+1}-1)\}_{i\geq 1}.$
The slope of $\overline{p_{i}p_{i+1}}$ is
$\frac{(2F_{2i+3}-1)-(2F_{2i+1}-1)}{(-F_{2i+2})-(-F_{2i})}$ $=-2\frac{F_{2i+2}}{F_{2i+1}}$
$=-2\left(\frac{\phi^{2i+2}-\psi^{2i+2}}{\phi^{2i+1}-\psi^{2i+2}}\right)$
$=-2\left(\phi+\frac{(\phi-\psi)(\psi^{2i+1})}{\phi^{2i+1}-\psi^{2i+1}}\right)$
$=-2\left(\phi+\frac{\psi^{2i+1}}{F_{2i+1}}\right).$
Because $\psi<0$, this expression decreases as $i$ increases. So the points $p_{i}$ are in convex position, making them the vertices of $\mathcal{P}.$
Let $\ell$ be the line $y=-2\phi x.$ By Binet’s Formula, the vertical distance between $p_{i}$ and $\ell$ is
$2\phi\left(\frac{\phi^{2i}-\psi^{2i}}{\phi-\psi}\right)-\left(\frac{2\phi^{2i+1}-2\psi^{2i+1}}{\phi-\psi}-1\right)=1-2\psi^{2i}.$
In particular, as $i$ increases, the distance between $P_{i}$ and $\ell$ increases and approaches $1$
Let $L$ be the strip defined by $-2\phi-1<y=-2\phi x$. We have concluded that $\mathcal{P}\subseteq L$. Since the vertical width of $L$ is 1, no two integer points in $L$ have the same $y$ coordinate; because the slope of $\ell$ is less than $-2$, the $y$-coordinates of any two lattice points in $\mathcal{P}$ differ by at least 2. In particular, the $y$-coordinates of the points $p_{i}$ are distinct from the $y$-coordinates of any other integer points that $\mathcal{P}$ contains. If $Y$ be the set of $y$-coordinates of non-vertex lattice points in $\mathcal{P}$. Then $A=\mathbb{Z}\setminus Y$ is a 2-syndetic set and $\mathcal{P}$ is an infinite empty polygon in $A\times A$. ∎
2.3 Arithmetic Congruence Sets
We begin directly with the proof.
Proof of Theorem 1.3.
Consider an empty polygon $\mathcal{P}$ with vertices in $A^{d}.$ Suppose for the sake of contradiction that $\mathcal{P}$ has two vertices $u$ and $v$ with $v-u\in m\mathbb{Z}^{d}$. For each $11\leq i\leq m-$ 1, let $c_{i}=u+\frac{i}{m}(v-u)$, which is an integer point. Fix $1\leq i\leq d$, and consider the $i$th coordinates of the points $c_{1},\dots,c_{m-1},$ which form an arithmetic progression modulo $m.$
  • $\circ$
    If this arithmetic progression is constant modulo $m,$ then since the $i$th coordinates of $u$ and $v$ are in $A,$ the $i$th coordinates of $c_{1},\dots,c_{m-1}$ are also in the set $A.$
  • $\circ$
    If this arithmetic progression is nonconstant, then any residue appears in the progression at most $\tfrac{m}{p}$ times. Therefore, the $i$th coordinates of $c_{1},\dots,c_{m-1}$ contain at most $\tfrac{m}{p}(m-k)$ residues not in $A.$
In either case, the $i$th coordinates of $c_{1},\dots,c_{m-1}$ contain at most $\tfrac{m}{p}(m-k)$ residues not in $A$. Summing over all $i,$ there are at most $\tfrac{m}{p}(m-k)d$ coordinates among $c_{1},\dots,c_{m-1}$ which are not in $A.$ However, we know $\tfrac{m}{p}(m-k)d<m-1.$ So by the Pigeonhole Principle, there is some $c_{i}$ with all of its coordinates in $A.$ This violates the emptiness condition.
Thus, there is at most one vertex of any given sequence of remainders modulo $m,$ so there are at most $k^{d}$ total vertices in $\mathcal{P}.$
It is not true in general that $h(A^{d})\leq|S|^{d}$. For example:
Proposition 2.4.
if $S=\{0,1\}$ and $A=S+3\mathbb{Z}$, then $h(A^{2})=8$.
Proof.
A construction of $8$ points is shown in Figure 2.
[…]
We claim that any empty polygon $P$ has at most two vertices whose modulo-$3$ residues are $(0,0)$. Then, by applying the symmetry of $A$, each of the four possible residues in $(\mathbb{Z}/3\mathbb{Z})^{2}$ can appear at most twice among the vertices of an empty polygon—so any empty polygon has at most $8$ vertices.
Suppose for the sake of contradiction that $(3x_{1},3y_{1}),(3x_{2},3y_{2}),(3x_{3},3y_{3})$ are all vertices of $P$. Since $P$ is empty, the centroid $(x_{1}+x_{2}+x_{3},y_{1}+y_{2}+y_{3})$ must have a coordinate that is congruent to $2$ (mod 3). We may assume $x_{1}+x_{2}+x_{3}\equiv 2\pmod{3}.$ Since the residues of $x_{1},x_{2},x_{3}$ take values in $\{0,1\}$, they cannot be distinct modulo $3$. The arguments are all the same, so assume $x_{1}\equiv x_{2}\pmod{3}.$ The congruence conditions guarantee that following convex combination are integer points:
$(2x_{1}+x_{2},2y_{1}+y_{2}),$
$(2x_{2}+x_{1},2y_{2}+y_{1}).$
Since the $x$-coordinates of both points are 0 (mod 3), the $y$-coordinates of both points must have remainder 2 (mod 3). But this is impossible, since the sum of the $y$-coordinates is $0\pmod{3}$. Therefore $P$ contains at most 2 vertices congruent to $(0,0)$ mod $3\mathbb{Z}^{2}$. ∎
3 Variations on Emptiness
3.1 Interior lattice points
In this section, we focus on lattice polytopes. An empty polytope contains no lattice point anywhere except its vertices. We now examine a less restrictive condition, which forbids lattice points only in the interior of a polytope. Polytopes satisfying this condition are called hollow.
In the spirit of Doignon’s Theorem, we could attempt to maximize the number of lattice points in the convex hull of a hollow polytope. However, hollow polytopes can contain arbitrarily many lattice points: Take, for example, a lattice rectangle $[0,1]\times[0,n]$. We could instead try to maximize the number of vertices, but that will not work either:
Theorem 3.1.
For any $d\geq 3$, there is convex polytope $\mathcal{P}$ in $\mathbb{Z}^{d}$ with arbitrarily many vertices such that the only lattice points in $\mathcal{P}$ are in its vertices, edges, or two-dimensional faces.
Proof.
Let $x_{1},\dots,x_{n}$ be the vertices of a convex $n$-gon in $\mathbb{Z}^{2}\subset\mathbb{Z}^{d}$, and let $e_{i}$ be the $i$th standard basis vector. Then $\operatorname{conv}(x_{1},\dots,x_{n},e_{3},\dots,e_{d})$ is a polytope with the desired property. ∎
In order to avoid such constructions, we must impose restrictions on the sets of points we are considering. Let $\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})$ denote the maximum number of vertices of a hollow lattice polytope whose vertices are in general position. (In this paper, a set is in “general position” if no $d+1$ points of the set are contained in a hyperplane.) It is not obvious that $\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})$ is finite, either; while the general position condition is restrictive, we still have the freedom of allowing many integer points on the boundary of the convex hull.
In fact, $\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})$ is finite, and one crucial fact implying this is that in a polytope whose vertex set is in general position, every facet is a simplex. Therefore, we also define $\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})$ as the maximum number of vertices in a hollow simplicial lattice polytope. Since any hollow lattice polytope with vertices in general position is simplicial, $\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})\leq\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})$. Our first result is that $\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})$ is finite.
Theorem 3.2.
$\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})\leq 2^{d}$.
Proof.
We will show that for any hollow simplicial polytope, we can construct an empty polytope with the same number of vertices.
Assume that $P$ is a simplicial lattice polytope whose boundary contains a lattice point $x$ that is not a vertex. Let $y$ be a vertex of the minimal face of $P$ that contains $x$, and let $Q$ be the polytope whose vertex set is $\{x\}\cup\operatorname{vert}(P)\setminus\{y\}$. Then $Q$ is a simplicial lattice polytope with strictly fewer non-vertex lattice points on its boundary and the same number of vertices as $P$. By induction, there is an empty polytope with the same number of vertices as $P$, so Doignon’s theorem implies that $\lvert\operatorname{vert}(P)\rvert\leq 2^{d}$. ∎
The bound in Theorem 3.2 is not necessarily sharp for $d>2$ because the facets of parallelepipeds (the equality cases of Doignon’s Theorem) are not all simplices. We do not have a better bound for simplicial polytopes. But for polytopes whose vertices are in general position, we can obtain a much better bound using the concept of lattice width.
Definition 3.3.
The width of a convex body $K\subseteq\mathbb{R}^{d}$ in the direction of a vector $v\in\mathbb{R}^{d}$ is
$w_{v}(K)=\max_{x\in K}\langle x,v\rangle-\min_{x\in K}\langle x,v\rangle.$
The lattice width of a convex body $K$ is
$w_{L}(K)=\min_{v\in\mathbb{Z}^{d}\setminus\{0\}}w_{v}(K).$
The lattice hyperplanes orthogonal to an integer vector $v$ are those hyperplanes orthogonal to $v$ that intersect $\mathbb{Z}^{d}$. Two adjacent lattice hyperplanes orthogonal to $v$ are distance $1/\|v\|$ apart. So if $w_{v}(K)=\alpha$, at most $\alpha+1$ lattice hyperplanes orthogonal to $v$ intersect $k$, and if $w_{L}(K)=\alpha$, then there is a lattice vector $v$ such that at most $\alpha+1$ lattice hyperplanes orthogonal to $v$ intersect $K$.
In 1948, Khinchine [14] proved the fundamental fact that any convex body containing no lattice points has bounded lattice width. The minimal upper bound on this lattice width is called the flatness constant. A series of works reduced Khinchine’s upper bound; Kannan and Lovász in 1988 provided an elegant proof that the flatness constant in $\mathbb{R}^{d}$ is at most $d^{2}$ [13]. On the other hand, the flatness constant is at least $d$: The interior of the lattice polytope $\operatorname{conv}(0,de_{1},\dots,de_{d})$ is a convex set with lattice width $d$ containing no lattice points. Recently, Reis and Rothvoss obtained an upper bound that is near optimal:
Theorem 3.4 (Flatness theorem, Reis–Rothvoss [16]).
If $K$ is a convex body in $\mathbb{R}^{d}$ that contains no lattice points, $w_{L}(K)=O\big{(}d(\log d)^{3}\big{)}$.
We can apply this result to obtain an improved upper bound for $\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})$.
Theorem 3.5.
$\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})=O\big{(}d^{2}(\log d)^{3}\big{)}$.
Proof.
Let $X$ be a set of integer points in $\mathbb{Z}^{d}$ in general position with $\operatorname{vert}\!\big{(}\!\operatorname{conv}(X)\big{)}\cap\mathbb{Z}^{d}=X$, and let $c\in\operatorname{conv}(X)$. Then $K:=\frac{1}{2}(\operatorname{conv}(X)-c)+c$ is a convex body that contains no lattice points. Using Theorem 3.4, we obtain a vector $v$ such that $K$ intersects at most $O(d(\log d)^{3})$ lattice hyperplanes orthogonal to $v$; therefore $\operatorname{conv}(X)$ intersects at most $O(d(\log d)^{3})$ lattice hyperplanes orthogonal to $v$, as well. Each lattice hyperplane contains at most $d$ points of $X$, so $X$ can contain no more than $O(d^{2}(\log d)^{3})$ points. ∎
As for lower bounds, the lattice simplex $\operatorname{conv}(0,e_{1},\dots,e_{d})$ shows that $\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})\geq\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})\geq d+1$. A simple construction slightly improves this lower bound.
Proposition 3.6.
$\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})\geq 2d$
Proof.
Let $X=\{e_{i}\}_{i=1}^{d}\cup\{e_{1}-e_{i}\}_{i=1}^{d}$, where $e_{i}$ is the standard unit basis vector in $\mathbb{R}^{d}$. Since $X$ is the union of points on the consecutive lattice hyperplanes $\{x\in\mathbb{Z}^{d}:\langle x,\mathbf{1}\rangle=0\}$ and $\{x\in\mathbb{Z}^{d}:\langle x,\mathbf{1}\rangle\}$, the lattice polytope $\operatorname{conv}(X)$ contains no lattice points in its interior. To prove that $\operatorname{conv}(X)$ is simplicial, let $u$ be an arbitrary nonzero vector in $\mathbb{R}^{d}$, and let $X_{u}$ be the set of points in $X$ that maximize the linear functional $\langle u,\cdot\rangle$. We need to show that $|X_{u}|\leq d$ for every $u\in\mathbb{R}^{d}\setminus\{0\}$.
We will prove that it is not possible that both $e_{i}$ and $e_{1}-e_{i}$ are in $X_{u}$, which immediately implies that $|X_{u}|\leq d$. To that end, suppose for contradiction that $e_{i},e_{1}-e_{i}\in X_{u}$. Then $u_{i}=u_{1}-u_{i}$. If $u_{1}=0$, then $0=u_{j}=\max_{x\in X}\langle x,u\rangle$, which is impossible (since $u\neq 0$). Otherwise:
  • $\circ$
    if $u_{1}>0$, then $u_{1}=2u_{j}>u_{j}$, which contradicts maximality;
  • $\circ$
    if $u_{1}<0$, then $\max_{x\in X}\langle x,u\rangle]=u_{1}<0$, which is clearly false, since $0\in X$.
So at most one of $e_{i}$ and $e_{1}-e_{i}$ can be in $X_{u}$, and $|X_{u}|\leq d$. ∎
3.2 Lattice segments
Next, we investigate polytopes which cannot have certain point sets in their convex hull. The simplest possible example of this is prohibiting segments of a given length.
Definition 3.7.
We use $[x,y]$ to denote the line segment with endpoints $x$ and $y$. If $x,y\in\mathbb{Z}^{d}$ with $|[x,y]\cap\mathbb{Z}^{d}|=k$, we say that $[x,y]$ is a lattice segment of length $k-1$.
Notice that a single point has length 00. We first consider the maximum number of lattice points a polytope can contain without containing a lattice segment of length $k$.
Proposition 3.8.
If $P$ is a convex lattice polytope in $\mathbb{R}^{d}$ and $P$ does not contain a lattice segment of length $k$, then $|P\cap\mathbb{Z}^{d}|\leq k^{d}$.
Proof.
If $|P\cap\mathbb{Z}^{d}|>k^{d}+1$, then there are two lattice points $x,y\in P\cap\mathbb{Z}^{d}$ whose coordinates are equal modulo $k$. If we set $z=\frac{1}{k}(y-x)\in\mathbb{Z}^{d}$, then $x,x+z,x+2z\dots,x+kz=y$ is a set of $k+1$ colinear lattice points contained inside $P$; so $P$ contains a lattice segment of length at least $k$. ∎
Since the hypercube $[1,k]^{d}$ contains exactly $k^{d}$ lattice points but no set of $k+1$ colinear lattice points, this inequality is tight. On the other hand, such lattice polytopes contain many fewer vertices.
Proposition 3.9.
Any convex lattice polytope in $\mathbb{R}^{d}$ that does not contain a lattice segment of length $k$ has $O_{d}(k^{d-1})$ vertices.
Proof.
We proceed by induction. The theorem is true if $d=1$, since every polytope in $\mathbb{R}^{1}$ has at most 2 vertices. Suppose that $P$ is a convex lattice polytope in $\mathbb{R}^{d}$ with no lattice segment of length $k$. If $|P\cap k\mathbb{Z}^{d}|\geq 2$, then $P$ contains a lattice segment of length $k$. Thus $|P\cap k\mathbb{Z}^{2}|\leq 1$. By translating $P$ by an integer vector, we may assume that $P\cap(2k)\mathbb{Z}^{2}=\emptyset$.
In this case, $\frac{1}{2k}P$ does not intersect the integer lattice. By the Flatness Theorem, there is a constant $c_{d}$ so that $\frac{1}{2k}P$ has lattice width at most $c_{d}$; so $P$ has lattice width at most $2c_{d}k$. This means that there is an integer vector $v$ such that at most $12c_{d}k+$ 1 lattice hyperplanes orthogonal to $v$ intersect $P$. Every vertex of $P$ lies on one of these hyperplanes, and each hyperplane has at most $O_{d}(k^{d-2})$ vertices by induction. So $P$ has at most $O_{d}(k^{d-1})$ vertices. ∎
On the other hand, there is a lattice polytope with almost the same number of vertices and no lattice segment of length $k$.
Proposition 3.10.
There is a convex lattice polytope in $\mathbb{R}^{d}$ with $\Theta_{d}(k^{(d-1)\frac{d}{d+1}})$ vertices and no lattice segment of length $k$.
Proof.
Take $P=\operatorname{conv}\big{(}\frac{k-1}{2}B^{d}\cap\mathbb{Z}^{d}\big{)}$, where $B^{d}$ is the unit ball in $\mathbb{R}^{d}$. Since $P\subset\frac{k-1}{2}B^{d}$, it does not contain a lattice segment of length $k$. On the other hand, Bárány [5] proved that $P$ has $\Theta_{d}(k^{(d-1)\frac{d}{d+1}})$ vertices. ∎
4 Open Problems
The main unsolved problem in Section 2 is bounding the number of vertices of empty polygons in exponential lattices in 3 or more dimensions. In particular, it is still unknown whether any three-dimensional exponential lattice contains an empty polygon with infinitely many vertices.
Problem 1.
Is $h\big{(}L_{3}(\alpha)\big{)}<\infty$?
While the majority of our proof of Theorem 1.1 does extend to higher dimensions, we cite Lemma 2.1 to handle a special case, and its proof in [1] relies on ordering the edge slopes of a polygon. Since facets in higher dimensions cannot be ordered in this way, a new method is required.
The growth of $h(L_{d}(\alpha))$—assuming, of course, it is finite—is also unknown. In the two-dimensional case, Ambrus et al. [1] note that if $\alpha=1+\tfrac{1}{x},$ the best known bounds in the plane are
$\lfloor\sqrt{x}\rfloor\leq h(L_{2}(\alpha))\leq O(x\log_{2}(x)).$
Problem 2.
Improve the asymptotic bounds on $h\big{(}L_{2}(1+\frac{1}{x})\big{)}$ as $x\to\infty$.
In higher dimensions, as $x\to\infty,$ Section 2.1.2 shows that
$\Omega(x^{(d-1)/2})\leq h(L_{d}(\alpha)),$
though of course we have proven no upper bound.
Our bounds on arithmetic congruence sets are strongest when $m$ is prime and $S$ is a significant proportion of $\{1,2,\dots,m\}$. What happens when these conditions aren’t true? There may be much to explore in this direction.
Problem 3.
Prove further bounds on Helly numbers for arithmetic congruence sets.
In Section 3, there are several additional open problems. First, although we obtain lower and upper bounds for $\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})$, they are far apart asymptotically: We only know that $2d\leq\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})\leq 2^{d}$. Is the growth rate of $\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})$ truly exponential, or is it much slower? There is an asymptotic difference in the bounds for $\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})$, as well, though the gap is less dramatic.
Problem 4.
Improve the asymptotic bounds for $\operatorname{hol}_{\textup{gp}}(\mathbb{Z}^{d})$ and $\operatorname{hol}_{\triangle}(\mathbb{Z}^{d})$.
Finally, our bounds for the number of vertices of a lattice polygon in $\mathbb{R}^{d}$ that does not contain a lattice segment of length $k$, although close, do not quite match. We suspect that our lower or upper bound is tight, but make no conjecture which it is.
Problem 5.
Establish tight asymptotic bounds on the maximum number of vertices a lattice polytope that does not contain a lattice segment of length $k$ can have.
5 Acknowledgements
This research was conducted under the auspices of the MIT PRIMES USA program. Dillon was further supported by a National Science Foundation Graduate Research Fellowship under Grant No. 2141064.